OpenStudy (anonymous):

Jerry had quarts of milk in his refrigerator. He used of the milk for a recipe. How much milk is left in the refrigerator? Express your answer in simplest form. A. quarts B. quarts C. quarts D. quarts

OpenStudy (anonymous):

@Miracrown

OpenStudy (anonymous):

@sbuck98

OpenStudy (photon336):

is it me or are the numbers missing @pink33

OpenStudy (anonymous):

ph whoops! yes

OpenStudy (anonymous):

Jerry had 3 3/4 he used 1/2 of the milk for a recipe. how much milk is left in the refrigerator?

OpenStudy (photon336):

So he had \[3\frac{ 3 }{ 4 }~quarts \] and he used\[\frac{ 1 }{ 2 }, of~the~milk \] for his recipe

OpenStudy (anonymous):

A. 1 1/ 2 B. 1 3/4 C.1 7/8 D. 2 3/4

OpenStudy (anonymous):

uh-huh

OpenStudy (photon336):

can you change 3 3/4 to an improper fraction?

OpenStudy (anonymous):

no

OpenStudy (photon336):

do you know how to do this?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

that's why i posted, lol

OpenStudy (photon336):

\[3\frac{ 3 }{ 4 }\] FIRST we multiply the whole number in this case 3 by the denominator do this and tell me what you get

OpenStudy (anonymous):

1 1/2

OpenStudy (photon336):

so the first thing we do is multiply the whole number which is 3 by the denominator which is 4 and we get 12

OpenStudy (photon336):

Then we add 12 to the numerator which is 3 and we get 15 \[\frac{ 15 }{ 4 } \]

OpenStudy (anonymous):

ok

OpenStudy (photon336):

we know that he used half of his milk for the recipe

OpenStudy (photon336):

so to represent this we would divide 15/4 by 2 do you know how to divide fractions? \[\frac{ 15 }{ 4 } \div 2 \]

OpenStudy (anonymous):

no

OpenStudy (photon336):

okay so first. we keep the first fraction, change the sign from division to multiplication and then flip the second fraction.

OpenStudy (photon336):

\[\frac{ 15 }{ 4 } * \frac{ 1 }{ 2 } = ? \]

OpenStudy (photon336):

so @pink33 what would be our answer

OpenStudy (photon336):

do you know how to multiply those two fractions?

OpenStudy (anonymous):

1 7/8

OpenStudy (photon336):

there you go. one more problem can you solve this? based on what we said \[\frac{ 1 }{ 2 } \div \frac{ 1 }{ 3 } = ? \]

OpenStudy (photon336):

and also can you change \[2\frac{ 2 }{ 5 }\] into an improper fraction. tag me when you have an answer @pink33

OpenStudy (anonymous):

1/6

OpenStudy (photon336):

remember it's division. do you recall the steps ?

OpenStudy (photon336):

Keep change flip

OpenStudy (anonymous):

is tht the answer?

OpenStudy (photon336):

so 1 7/8 was the answer to your question but the other one was incorrect. \[\frac{ 1 }{ 2 } \div \frac{ 1 }{ 3 } = \frac{ 1 }{ 2 } * \frac{ 3 }{1 } = \frac{ 3 }{ 2 }\]

OpenStudy (photon336):

take a look at the procedure then do this one \[\frac{ 1 }{ 4 } \div \frac{ 1 }{ 6 } = ? \]

OpenStudy (anonymous):

Joel took of an hour to do his chores. It took Halley the time it took Joel to do his chores. How much time did it take Halley to do her chores? Express your answer in simplest form. A. hr B. hr C. hr D. hr

OpenStudy (anonymous):

Joel took 4/5 of an hour it took hailey 1/8

OpenStudy (anonymous):

@Photon336

OpenStudy (anonymous):

A. 1/20 hr B. 1/10 hr C. 4/13 D. 5/32

OpenStudy (photon336):

what do you think based on what we discussed?

OpenStudy (anonymous):

idk

OpenStudy (photon336):

at least try it

OpenStudy (anonymous):

ok it a

OpenStudy (photon336):

how did you get that?